OpenStudy (anonymous):

what is the value of k if point S(-4,k) and T(-1,0) are sqrt 73 ts apart

OpenStudy (anonymous):

if THERE Are two points A (x,y) and B(m,n).. the distance between them is \[\sqrt{(x-m)^2+(y-n)^2}\]

OpenStudy (anonymous):

you know the distance is 73, just put the values in the equation.

OpenStudy (anonymous):

??/

OpenStudy (anonymous):

\[\sqrt{(-4+1)^2+(k-0)^2}=\sqrt{73}\]

OpenStudy (anonymous):

now, can you do it?

OpenStudy (anonymous):

mmm imma try and get back at you thanks alot:)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

i probably did it so wrong but is it 0+k2^+148

OpenStudy (anonymous):

well,you square both sides .. you will get \[{(-3)^2+k^2}=73\]

OpenStudy (anonymous):

\[so, 9+k^2=73\]

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

\[k^2=64 \]

OpenStudy (anonymous):

so, k=8

OpenStudy (anonymous):

\[\pm8\]